# : Generalizing Large Language Models for E-commerce

### v1

@@ 5 Experimental Setup
| General-purpose LLMs |
(1) GPT-4 TurboΒ (OpenAI, 2023), (2) Gemini ProΒ (Team etΒ al., 2023), (3) Claude 2.1Β (Anthropic, 2023), (4) Llama-2 13B-chatΒ (Touvron etΒ al., 2023b), (5) Mistral-7B Instruct-v0.2Β (Jiang etΒ al., 2023) |
E-commerce LLM |
(1) EcomGPTΒ (Li etΒ al., 2023b): To the best of our knowledge, it is the only open-source e-commerce LLM. |
@@
| (SectionΒ A.1.1) |
(1) SUOpenTagΒ (Xu etΒ al., 2019), (2)AVEQAΒ (Wang etΒ al., 2020): Both methods scan the product information, token by token, to extract the values of the specified attributes. |
| (SectionΒ A.1.2) |
(1) RGCNΒ (Schlichtkrull etΒ al., 2018): It utilizes a graph convolutional network to capture relations between products. (2) DeBERTaV3Β (He etΒ al., 2022): It predicts product relations from product titles. |
| (SectionΒ A.2.1) |
(1) BERTweetΒ (Nguyen etΒ al., 2020): It is a pre-trained language model for English tweets, and achieves superior performance over RoBERTaΒ (Liu etΒ al., 2019) and XLM-RΒ (Conneau etΒ al., 2019) on . (2) P5Β (Geng etΒ al., 2022): It is a pre-trained LLM for from e-commerce user reviews. |
| (SectionΒ A.2.2) |
(1) gSASRecΒ (Petrov & Macdonald, 2023), (2) RecformerΒ (Li etΒ al., 2023): Both methods leverage TransformerΒ (Vaswani etΒ al., 2017) to predict the next product of users' interest based on users' historical activities on products. |
| (SectionΒ A.4.2) |
(1) GPT-4 TurboΒ (OpenAI, 2023): To the best of our knowledge, there are no models specifically designed for . Thus, GPT-4 Turbo is used as the SoTA task-specific model in this task as it has outstanding performance in general question answeringΒ (OpenAI, 2023). |
@@
We compare against 3 categories of baseline models: (1) general-purpose LLMs, (2) e-commerce LLMs, and (3) SoTA task-specific models. TableΒ 2 lists all the baseline models. Note that for the task-specific models, we only use the best (i.e., SoTA) models for each individual task based on literature, so as to enable schematic comparison between generalist modeling from and the task-specific modeling. We conduct IND and OOD tests on respective test datasets (SectionΒ 3.4) for all the models. TableΒ A2 lists the evaluations conducted for different tasks. Details on using/training baseline models are in AppendixΒ C. The prompt templates used in the evaluations are in AppendixΒ D.
@@
We perform 1-shot evaluation on all the general-purpose LLMs. Existing workΒ (Li etΒ al., 2023b) tests LLMs on e-commerce tasks in a 0-shot setting. Meanwhile, in-context examples can notably benefit LLMsΒ (Brown etΒ al., 2020). Thus, in our case, we conduct 1-shot evaluation, balancing the computing cost and model performance, to enable stronger performance from the general-purpose LLMs. Although extensive prompt engineering and many in-context examples could enable better performance from LLMs, it is less practical in real e-commerce applications; for example, asking users to provide many in-context examples to the e-commerce platform, if ever feasible, can reduce user engagement. Also, it is not cost- and energy-efficient to do large-scale, few-shot in-context learning on LLMs over large test sets like 's.
E-commerce LLMsΒ Β For EcomGPT, we use the checkpoint released by its authors. We conduct both 0-shot and 1-shot evaluations on EcomGPT, as we empirically observe that 1-shot may result in better performance than 0-shot for EcomGPT (EcomGPT originally conducts 0-shot evaluation). Therefore, we report the best performance of the two evaluations on each task.
@@
| Β Β Β GPT-4 Turbo | 0.495 | 0.326 | 0.753 | 0.516 | 0.387 | 0.611 | 0.195 | 0.875 | 0.649 | 0.858 |
| Β Β Β Gemini Pro | 0.396 | 0.136 | 0.867 | 0.470 | 0.269 | 0.584 | 0.248 | 0.821 | 0.506 | 0.855 |
| Β Β Β Claude 2.1 | 0.381 | 0.275 | 0.523 | 0.415 | 0.066 | 0.655 | 0.273 | 0.821 | 0.280 | 0.841 |
| Β Β Β Llama-2 13B-chat | 0.002 | 0.333 | 0.434 | 0.188 | 0.056 | 0.504 | 0.252 | 0.815 | 0.623 | 0.811 |
| Β Β Β Mistral-7B Instruct-v0.2 | 0.369 | 0.324 | 0.613 | 0.470 | 0.164 | 0.529 | 0.305 | 0.842 | 0.588 | 0.853 |
| Β Β Β EcomGPT | 0.000 | 0.091 | 0.648 | 0.188 | 0.042 | 0.540 | 0.170 | 0.000 | 0.086 | 0.669 |
| Β Β Β SoTA task-specific model | 0.546 | 0.588 | 0.995 | 0.573 | 0.265 | 0.703 | 0.389 | 0.859 | 0.830 | 0.858 |
| Β Β Β | 0.582 | 0.611 | 0.995 | 0.648 | 0.526 | 0.684 | 0.501 | 0.870 | 0.851 | 0.841 |
| Β Β Β | 0.662 | 0.558 | 0.995 | 0.639 | 0.542 | 0.696 | 0.305 | 0.876 | 0.846 | 0.842 |
| Β Β Β | 0.509 | 0.518 | 0.991 | 0.596 | 0.479 | 0.650 | 0.392 | 0.870 | 0.846 | 0.842 |
| Β Β Β improvement (%, avg: 10.7) | 21.2 | 3.9 | 0.0 | 13.1 | 40.1 | -1.0 | 28.8 | 0.1 | 2.5 | -1.9 |
@@
In this table, ``F1*", ``Macro F1", ``F1", ``HR@1", ``Accuracy", ``NDCG" and ``F" are the primary evaluation metrics in respective tasks (AppendixΒ A). For each task, the best baseline performance is underlined, and the overall best performance is in bold. The row ``improvementβ presents the percentage improvement of the best-performing model over the best-performing baseline model (underlined) in each task. We also include the average (`avg') improvement across all the tasks in the table.
@@ 6 Experimental Results
We evaluate the models using the test sets of individual tasks. For each task, we employ multiple metrics for a comprehensive evaluation (AppendixΒ A). For the sake of clarity, in this section, we present the performance only in terms of the primary evaluation metric for each task (TableΒ A1). As discussed in AppendixΒ E, Llama-2 13B-chat, Mistral-7B Instruct-v0.2, and Phi-2 are the best base models for , , and , respectively. Thus, by default, , , and refer to those tuned from these base models, respectively.
Main ResultsΒ Β Our comprehensive experiments yield the following main results: (1) models demonstrate the best performance on almost all the IND tasks, with a significant average improvement of 10.7% over the general-purpose LLMs, e-commerce LLMs, and the SoTA task-specific models across the 10 tasks (SectionΒ 6.1). (2) models show outstanding generalizability to OOD products and surpass the best baselines with a remarkable average improvement of 9.3% in OOD evaluation (SectionΒ 6.2). (3) By training over diverse instructions, is equipped with strong generalizability to unseen instructions (SectionΒ 6.3). (4) Trained on all the tasks in ECInstruct together, exhibits similar or better performance than models trained on each individual task (SectionΒ 6.4). (5) models benefit from larger instruction training data for e-commerce tasks (SectionΒ 6.5).
@@ 6.1 In-domain Evaluation
TableΒ 3 shows the performance of and all the baseline methods in IND evaluation, where models are trained using training set (i.e., including all the tasks), and the SoTA task-specific models are trained using the task-specific training data. Among the SoTA task-specific models, we report the results of only the best-performing model on each task. The complete results on each task are presented in AppendixΒ F.
| β Model | ||||||
| F1* | M-F1 | M-F1 | HR@1 | F1 | F | |
| β GPT-4 Turbo | 0.397 | 0.392 | 0.510 | 0.198 | 0.680 | 0.860 |
| β Gemini Pro | 0.275 | 0.123 | 0.454 | 0.116 | 0.552 | 0.856 |
| β Claude 2.1 | 0.410 | 0.277 | 0.369 | 0.036 | 0.245 | 0.842 |
| β Llama-2 13B-chat | 0.000 | 0.324 | 0.178 | 0.050 | 0.644 | 0.808 |
| β Mistral-7B | 0.264 | 0.327 | 0.438 | 0.108 | 0.608 | 0.851 |
| β Instruct-v0.2 | ||||||
| β EcomGPT | 0.001 | 0.096 | 0.178 | 0.023 | 0.140 | 0.722 |
| β SoTA | 0.269 | 0.507 | 0.567 | 0.081 | 0.853 | 0.860 |
| β task-specific model | ||||||
| β | 0.335 | 0.558 | 0.629 | 0.273 | 0.867 | 0.841 |
| β | 0.367 | 0.502 | 0.640 | 0.280 | 0.878 | 0.840 |
| β | 0.302 | 0.520 | 0.565 | 0.241 | 0.879 | 0.840 |
| β improvement | -10.5 | 10.1 | 14.1 | 41.4 | 3.0 | -2.2 |
| β (%, avg: 9.3) |
β’
In this table, ``M-F1" represents macro F1. The columns in this table have the same meanings as those in TableΒ 3.
@@
Comparison between and E-commerce LLMsΒ Β According to TableΒ 3, models demonstrate substantial improvement over the existing e-commerce LLM EcomGPT on all the tasks. For example, surpasses EcomGPT remarkably by 244.6% on , and by 53.5%. Both and EcomGPT are instruction-tuned LLMs for e-commerce. However, our instruction dataset for fundamentally differs from EcomGPT's EcomInstruct dataset: includes only real-world data and real-world e-commerce tasks, while EcomInstruct incorporate a considerable amount of synthetic data and tasks, which hinders its applicability in real e-commerce tasks.
@@ 6.2 Out-of-domain Evaluation
TableΒ 4 shows the performance of and baselines in OOD evaluation. Overall, we observe a similar trend to that in IND evaluation. Specifically, outperforms the SoTA task-specific models on OOD evaluation by a wide margin on 5 out of the 6 tasks, except for task, on which is comparable with the SoTA task-specific model (e.g., 2.2% difference). Similarly, outperforms the general-purpose LLMs on 4 out of the 6 tasks (i.e., , , and ). Across the 6 tasks, demonstrates a substantial average improvement of 18.9% over GPT-4 Turbo. Compared to EcomGPT, again shows terrific advantages on all the tasks in OOD evaluation. As shown in the literatureΒ (Lika etΒ al., 2014), cold start for new products has been an unsolved issue but a key driver in e-commerce applications. The OOD generalizability of to new products as demonstrated in TableΒ 4 makes a highly viable tool for e-commerce applications.
Note that GPT-4 and Claude 2.1 exhibit robust performance on and , indicating that these general-purpose LLMs are inherent with rich knowledge for solving extraction and question answering problems. However, they may still lack comprehensive knowledge to effectively perform multiple, diverse e-commerce tasks, such as on the other 4 tasks.
@@ 6.3 Generalizability to Unseen Instructions
In this table, ``single" and ``diverse" indicate that the models are tuned over single and diverse instructions, respectively. The best performance of each model (e.g., , abd ) is in bold, if the performance difference between the model tuned over single and diverse instructions exceeds 1%.
| Model | Training | ||||||||||
@@
In this table, ``Task-specific" indicates that the models are tuned on individual tasks; ``Generalist" represents tuning models using all tasks together. The best performance of generalist and task-specific models on each task is in bold, if the performance difference between the generalist and task-specific model exceeds 1%.
@@
These results illustrate the pivotal role of diverse instructions for e-commerce tasks, and underscore the utility of in generalizing LLMs to new instructions in e-commerce applications. The results also showcase the generalizability of to unseen instructions. As new users may apply instructions unseen in existing e-commerce data, such generalizability also indicates the potential of for cold-start users in e-commerce. It is noticeable that is able to generalize to unseen instructions (TableΒ 5) with a similar performance as it has on ``seen" instructions (TableΒ 3). For example, on , shows only a 1.4% difference in its performance in the former setting and in the latter setting (0.639 vs 0.648). This further demonstrates the strong generalizability of .
@@ 6.4 Generalist vs Task-specific Models
We compare the models tuned using all tasks (generalist models) with those tuned on individual tasks (task-specific models), and present the performance comparison in IND evaluation in TableΒ 6. The results of OOD evaluation are presented in AppendixΒ G.3. As shown in TableΒ 6, generalist demonstrates slightly stronger performance compared to task-specific on each task (e.g., 2.7% average improvement of generalist over all tasks except ). For example, on , generalist is comparable to the task-specific (0.582 vs 0.599). On , generalist demonstrates a substantial improvement of 17.3% compared to task-specific (0.611 vs 0.521). As evidenced by the results, by training on all tasks together, models enjoy strong versatility and could enable knowledge transfer across tasks for improved performance.

### v2

@@ 5 Experimental Setup
| General-purpose LLMs | (1) GPT-4 TurboΒ (OpenAI, 2023), (2) Gemini ProΒ (Team etΒ al., 2023), (3) Claude 2.1Β (Anthropic, 2023), (4) Llama-2 13B-chatΒ (Touvron etΒ al., 2023b), (5) Mistral-7B Instruct-v0.2Β (Jiang etΒ al., 2023) | |
| E-commerce LLM | (1) EcomGPTΒ (Li etΒ al., 2024): To the best of our knowledge, it is the only open-source e-commerce LLM. | |
@@
| (SectionΒ A.1.1) | (1) SUOpenTagΒ (Xu etΒ al., 2019), (2)AVEQAΒ (Wang etΒ al., 2020): Both methods scan the product information, token by token, to extract the values of the specified attributes. | |
| (SectionΒ A.1.2) | (1) RGCNΒ (Schlichtkrull etΒ al., 2018): It utilizes a graph convolutional network to capture relations between products. (2) DeBERTaV3Β (He etΒ al., 2022): It predicts product relations from product titles. | |
| (SectionΒ A.2.1) | (1) BERTweetΒ (Nguyen etΒ al., 2020): It is a pre-trained language model for English tweets, and achieves superior performance over RoBERTaΒ (Liu etΒ al., 2019) and XLM-RΒ (Conneau etΒ al., 2019) on . (2) P5Β (Geng etΒ al., 2022): It is a pre-trained LLM for from e-commerce user reviews. | |
| (SectionΒ A.2.2) | (1) gSASRecΒ (Petrov & Macdonald, 2023), (2) RecformerΒ (Li etΒ al., 2023): Both methods leverage TransformerΒ (Vaswani etΒ al., 2017) to predict the next product of usersβ interest based on usersβ historical activities on products. | |
| (SectionΒ A.4.2) | (1) GPT-4 TurboΒ (OpenAI, 2023): To the best of our knowledge, there are no models specifically designed for . Thus, GPT-4 Turbo is used as the SoTA task-specific model in this task as it has outstanding performance in general question answeringΒ (OpenAI, 2023). | |
@@
from the textual product information (e.g., titles). |
We compare against 3 categories of baseline models: (1) general-purpose LLMs, (2) e-commerce LLMs, and (3) SoTA task-specific models. TableΒ 2 lists all the baseline models. Note that for the task-specific models, we only use the best (i.e., SoTA) models for each individual task based on literature, so as to enable schematic comparison between generalist modeling from and the task-specific modeling. We conduct IND and OOD tests on respective test datasets (SectionΒ 3.4) for all the models. TableΒ A3 lists the evaluations conducted for different tasks. Details on using/training baseline models are in AppendixΒ D. The prompt templates used in the evaluations are in AppendixΒ E.
@@
We perform 1-shot evaluations on all the general-purpose LLMs. Existing workΒ (Li etΒ al., 2024) tests LLMs on e-commerce tasks in a 0-shot setting. Meanwhile, in-context examples can notably benefit LLMsΒ (Brown etΒ al., 2020). Thus, in our case, we conduct 1-shot evaluations, balancing the computing cost and model performance, to enable stronger performance from the general-purpose LLMs. Although extensive prompt engineering and many in-context examples could enable better performance from LLMs, it is less practical in real e-commerce applications; for example, asking users to provide many in-context examples of the e-commerce platform, if ever feasible, can reduce user engagement. Also, it is not cost- and energy-efficient to do large-scale, few-shot in-context learning on LLMs over large test sets like βs.
E-commerce LLMsΒ Β For EcomGPT, we use the checkpoint released by its authors. We conduct both 0-shot and 1-shot evaluations on EcomGPT, as we empirically observe that 1-shot may result in better performance than 0-shot for EcomGPT (EcomGPT originally conducts 0-shot evaluations). Therefore, we report the best performance of the two evaluations on each task.
@@
| GPT-4 Turbo | 0.495 | 0.326 | 0.753 | 0.516 | 0.387 | 0.611 | 0.195 | 0.875 | 0.649 | 0.858 |
| Gemini Pro | 0.396 | 0.136 | 0.867 | 0.470 | 0.269 | 0.584 | 0.248 | 0.821 | 0.506 | 0.855 |
| Claude 2.1 | 0.381 | 0.275 | 0.523 | 0.415 | 0.066 | 0.655 | 0.273 | 0.821 | 0.280 | 0.841 |
| Llama-2 13B-chat | 0.002 | 0.333 | 0.434 | 0.188 | 0.056 | 0.504 | 0.252 | 0.815 | 0.623 | 0.811 |
| Mistral-7B Instruct-v0.2 | 0.369 | 0.324 | 0.613 | 0.470 | 0.164 | 0.529 | 0.305 | 0.842 | 0.588 | 0.853 |
| EcomGPT | 0.000 | 0.091 | 0.648 | 0.188 | 0.042 | 0.540 | 0.170 | 0.000 | 0.086 | 0.669 |
| SoTA task-specific model | 0.546 | 0.588 | 0.995 | 0.573 | 0.265 | 0.703 | 0.389 | 0.859 | 0.830 | 0.858 |
| 0.582 | 0.611 | 0.995 | 0.648 | 0.526 | 0.684 | 0.501 | 0.870 | 0.851 | 0.841 | |
| 0.662 | 0.558 | 0.995 | 0.639 | 0.542 | 0.696 | 0.305 | 0.876 | 0.846 | 0.842 | |
| 0.509 | 0.518 | 0.991 | 0.596 | 0.479 | 0.650 | 0.392 | 0.870 | 0.846 | 0.842 | |
| improvement (%, avg: 10.7) | 21.2 | 3.9 | 0.0 | 13.1 | 40.1 | -1.0 | 28.8 | 0.1 | 2.5 | -1.9 |
@@
In this table, βF1*β, βMacro F1β, βF1β, βHR@1β, βAccuracyβ, βNDCGβ and βFβ are the primary evaluation metrics in respective tasks (AppendixΒ A). For each task, the best baseline performance is underlined, and the overall best performance is in bold. The row βimprovementβ presents the percentage improvement of the best-performing model over the best-performing baseline model (underlined) in each task. We also include the average (βavgβ) improvement across all the tasks in the table.
@@ 6 Experimental Results
We evaluate the models using the test sets of individual tasks. For each task, we employ multiple metrics for a comprehensive evaluation (AppendixΒ A). For the sake of clarity, in this section, we present the performance only in terms of the primary evaluation metric for each task (TableΒ A1). As discussed in AppendixΒ F, Llama-2 13B-chat, Mistral-7B Instruct-v0.2, and Phi-2 are the best base models for , , and , respectively. Thus, by default, , , and refer to those tuned from these base models, respectively.
Main ResultsΒ Β Our comprehensive experiments yield the following main results: (1) models demonstrate the best performance on almost all the IND tasks, with a significant average improvement of 10.7% over the general-purpose LLMs, e-commerce LLMs, and the SoTA task-specific models across the 10 tasks (SectionΒ 6.1). (2) models show outstanding generalizability to OOD products and surpass the best baselines with a remarkable average improvement of 9.3% in OOD evaluation (SectionΒ 6.2). (3) By training over diverse instructions, is equipped with strong generalizability to unseen instructions (SectionΒ 6.3). (4) Trained on all the tasks in together, exhibits similar or better performance than models trained on each individual task (SectionΒ 6.4). (5) models benefit from larger instruction training data for e-commerce tasks (SectionΒ 6.5).
@@ 6.1 In-domain Evaluation
TableΒ 3 shows the performance of and all the baseline methods in IND evaluation, where models are trained using training set (i.e., including all the tasks), and the SoTA task-specific models are trained using the task-specific training data. Among the SoTA task-specific models, we report the results of only the best-performing model on each task. The complete results of each task are presented in AppendixΒ G.
@@
Comparison between and E-commerce LLMsΒ Β According to TableΒ 3, models demonstrate substantial improvement over the existing e-commerce LLM EcomGPT on all the tasks. For example, surpasses EcomGPT remarkably by 244.6% on , and by 53.5%. Both and EcomGPT are instruction-tuned LLMs for e-commerce. However, our instruction dataset for fundamentally differs from EcomGPTβs EcomInstruct dataset: includes only real-world data and real-world e-commerce tasks, while EcomInstruct incorporates a considerable amount of synthetic data and tasks, which hinders its applicability in real e-commerce tasks.
@@ 6.2 Out-of-domain Evaluation
| Model | ||||||
| F1* | M-F1 | M-F1 | HR@1 | F1 | F | |
| GPT-4 Turbo | 0.397 | 0.392 | 0.510 | 0.198 | 0.680 | 0.860 |
| Gemini Pro | 0.275 | 0.123 | 0.454 | 0.116 | 0.552 | 0.856 |
| Claude 2.1 | 0.410 | 0.277 | 0.369 | 0.036 | 0.245 | 0.842 |
| Llama-2 13B-chat | 0.000 | 0.324 | 0.178 | 0.050 | 0.644 | 0.808 |
| Mistral-7B | 0.264 | 0.327 | 0.438 | 0.108 | 0.608 | 0.851 |
| Instruct-v0.2 | ||||||
| EcomGPT | 0.001 | 0.096 | 0.178 | 0.023 | 0.140 | 0.722 |
| SoTA | 0.269 | 0.507 | 0.567 | 0.081 | 0.853 | 0.860 |
| task-specific model | ||||||
| 0.335 | 0.558 | 0.629 | 0.273 | 0.867 | 0.841 | |
| 0.367 | 0.502 | 0.640 | 0.280 | 0.878 | 0.840 | |
| 0.302 | 0.520 | 0.565 | 0.241 | 0.879 | 0.840 | |
| improvement | -10.5 | 10.1 | 14.1 | 41.4 | 3.0 | -2.2 |
| (%, avg: 9.3) |
β’
In this table, βM-F1β represents macro F1. The columns in this table have the same meanings as those in TableΒ 3.
TableΒ 4 shows the performance of and baselines in OOD evaluation. Overall, we observe a similar trend to that in IND evaluation. Specifically, outperforms the SoTA task-specific models on OOD evaluation by a wide margin on 5 out of the 6 tasks, except for task, on which is comparable with the SoTA task-specific model (e.g., 2.2% difference). Similarly, outperforms the general-purpose LLMs on 4 out of the 6 tasks (i.e., , , , and ). Across the 6 tasks, demonstrates a substantial average improvement of 18.9% over GPT-4 Turbo. Compared to EcomGPT, again shows terrific advantages on all the tasks in OOD evaluation. As shown in the literatureΒ (Lika etΒ al., 2014), cold start for new products has been an unsolved issue but a key driver in e-commerce applications. The OOD generalizability of to new products as demonstrated in TableΒ 4 makes a highly viable tool for e-commerce applications.
Note that GPT-4 and Claude 2.1 exhibit robust performance on and , indicating that these general-purpose LLMs are inherent with rich knowledge for solving extraction and question-answering problems. However, they may still lack comprehensive knowledge to effectively perform multiple, diverse e-commerce tasks, such as on the other 4 tasks.
@@ 6.3 Generalizability to Unseen Instructions
In this table, βsingleβ and βdiverseβ indicate that the models are tuned over single and diverse instructions, respectively. The best performance of each model (e.g., , abd ) is in bold, if the performance difference between the model tuned over single and diverse instructions exceeds 1%.
| Model | Training Tasks | ||||||||||
@@
In this table, βTask-specificβ indicates that the models are tuned on individual tasks; βGeneralistβ represents tuning models using all tasks together. The best performance of generalist and task-specific models on each task is in bold, if the performance difference between the generalist and task-specific model exceeds 1%.
@@
These results illustrate the pivotal role of diverse instructions for e-commerce tasks, and underscore the utility of in generalizing LLMs to new instructions in e-commerce applications. The results also showcase the generalizability of to unseen instructions. As new users may apply instructions unseen in existing e-commerce data, such generalizability also indicates the potential of for cold-start users in e-commerce. It is noticeable that is able to generalize to unseen instructions (TableΒ 5) with a similar performance as it has on βseenβ instructions (TableΒ 3). For example, on , shows only a 1.4% difference in its performance in the former setting and in the latter setting (0.639 vs 0.648). This further demonstrates the strong generalizability of .
@@ 6.4 Generalist vs Task-specific Models
We compare the models tuned using all tasks (generalist models) with those tuned on individual tasks (task-specific models) and present the performance comparison in IND evaluation in TableΒ 6. The results of the OOD evaluation are presented in AppendixΒ H.3. As shown in TableΒ 6, generalist demonstrates slightly stronger performance compared to task-specific on each task (e.g., 2.7% average improvement of generalist over all tasks except ). For example, on , generalist is comparable to the task-specific (0.582 vs 0.599). On , generalist demonstrates a substantial improvement of 17.3% compared to task-specific (0.611 vs 0.521). As evidenced by the results, by training on all tasks together, models enjoy strong versatility and could enable knowledge transfer across tasks for improved performance.